1. Pyridine (C5H5N) is a particularly foul-smelling substance used in manufacturing pesticides and plastic resins. Calculate the pH of a 0.125 M solution of pyridine (Kb = 1.4×10-9 ).
1
Expert's answer
2020-04-13T02:24:55-0400

C5H5N + H2O <-> C5H5NH+ + OH-

"Kb = {\frac {[C_5H_5NH^+][OH^-]} {[C_5H_5N]}}=1.4*10^{-9}"

If x moles from 0.125 have dissociated then

"Kb = {\frac {[C_5H_5NH^+][OH^-]} {[C_5H_5N]}}={\frac {x^2} {0.125-x}}=1.4*10^{-9}"

0.125*1.4*10-9 - 1.4*10-9x = x2

x2 + 1.4*10-9x - 1.75*10-10 = 0

Solving the equation and getting the positive root we get

x = 1.323*10-5 = [OH-]

pH = 14 - pOH

pOH = -lg[OH-] = -lg(1.323*10-5) = 4.88

pH = 14 - 4.88 = 9.12

pH = 9.12